OpenStudy (anonymous):

Find the value of 10sqrt2 divided by sqrt5

OpenStudy (anonymous):

I think its 6.32, but not quite sure. Is this right anyone?

OpenStudy (anonymous):

The answers for the question are : A: 2 B: 2sqrt5 C: 2sqrt10 D:2/5sqrt10

OpenStudy (anonymous):

oh then I don't know. Sorry

jimthompson5910 (jim_thompson5910):

The idea is to rationalize the denominator by multiplying both top and bottom by \(\Large \sqrt{5}\) \[\Large \frac{10\sqrt{2}}{\sqrt{5}}\] \[\Large \frac{10\sqrt{2}*\sqrt{5}}{\sqrt{5}*\sqrt{5}}\] \[\Large \frac{10\sqrt{2*5}}{\sqrt{5*5}}\] \[\Large \frac{10\sqrt{10}}{\sqrt{25}}\] \[\Large \frac{10\sqrt{10}}{\sqrt{5^2}}\] \[\Large \frac{10\sqrt{10}}{5}\] \[\Large \frac{5*2\sqrt{10}}{5}\] \[\Large \frac{1*2\sqrt{10}}{1}\] \[\Large 2\sqrt{10}\] So \[\Large \frac{10\sqrt{2}}{\sqrt{5}}=2\sqrt{10}\]

OpenStudy (anonymous):

How did you go from 5*2sqrt10/5 to 1*2sqrt10/1?? Sorry, really bad at math...

jimthompson5910 (jim_thompson5910):

That's ok

jimthompson5910 (jim_thompson5910):

I canceled out the two '5' terms like so \[\Large \frac{5*2\sqrt{10}}{5}\] \[\Large \frac{\cancel{5}*2\sqrt{10}}{\cancel{5}}\] \[\Large \frac{1*2\sqrt{10}}{1}\]

jimthompson5910 (jim_thompson5910):

They cancel out since 5/5 = 1

OpenStudy (anonymous):

Ohhh, okayy Now I understand. Thank you :)

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

Wait I just want to make sure I'm right on this next one... 4sqrt(8)/aqrt(2) = 8... right?

jimthompson5910 (jim_thompson5910):

you are correct

jimthompson5910 (jim_thompson5910):

nice work

OpenStudy (anonymous):

Yay! Thank you! :)

jimthompson5910 (jim_thompson5910):

np